# Edge-Cloud Collaborative Computing on Distributed Intelligence and Model Optimization: A Survey

### v4

@@ III-C Resource-Aware Model Optimization
Resource-aware model optimization represents a critical advancement in ECCC systems, focusing on dynamic adaptation to heterogeneous and constrained computational environments. Advanced optimization strategies encompass three primary approaches: dynamic model selection based on real-time resource availability and application requirements, model splitting and offloading that strategically partition components across edge-cloud infrastructure, and reinforcement learning-based optimization for adaptive resource allocation. Dynamic model selection considers computational constraints, communication bandwidth, and energy limitations while maintaining QoS requirements, with techniquess ranging from online learning approaches to trust-aware selection strategies [116, 62]. Model partitioning leverages heterogeneous system capabilities by offloading computationally intensive components to resource-rich cloud while executing latency-sensitive parts at the edge, with sophisticated algorithms like MADDPG enhancing efficiency through workload-aware adaptation [23, 117]. Reinforcement learning mechanisms optimize resource allocation through environmental feedback, employing techniques like deep Q-learning and APPO to balance computational efficiency with system performance, achieving improved stability and reduced failure rates in IoT monitoring applications [21, 118]. Consequently, integrated optimization strategies establish a versatile framework for diverse domains including IoT networks, vehicular computing, and 5G environments, addressing the fundamental challenges of resource heterogeneity and dynamic operational constraints in distributed computing systems. For comprehensive technical details, evaluation methodologies, and implementation strategies, please see LABEL:sec:supp-tech-comparison of the Supplementary Material.
@@ III-D Neural Architecture Search
Neural architecture search (NAS) automates the design and optimization of neural network architectures to discover optimal configurations for edge-cloud collaborative environments, systematically exploring architectural configurations tailored to distributed computing constraints. Advanced automated methodologies encompass three primary paradigms: reinforcement learning-based NAS where agents learn optimal architecture selection through performance feedback and reward systems that incorporate edge-cloud deployment constraints; evolutionary algorithm-based NAS that employs natural selection principles to evolve architecture populations through iterative selection, mutation, and crossover operations; and gradient-based NAS that formulates architecture search as continuous optimization problems enabling efficient exploration through established gradient descent techniques [119, 120, 22]. Reinforcement learning approaches demonstrate particular effectiveness in resource-constrained environments by directly incorporating computational, memory, and energy constraints into reward functions, enabling discovery of architectures specifically optimized for edge device limitations [25, 121]. Evolutionary algorithms provide exceptional adaptability to dynamic edge-cloud environments through inherent population diversity and exploration capabilities, with applications ranging from scientific workflow optimization to adaptive resource allocation [122, 123]. Gradient-based methods achieve superior search efficiency by representing architecture spaces differentiably, reducing search time significantly while enabling exploration of complex architectural configurations, though challenges remain in handling operation correlations and incorporating device constraints directly into the search process [124, 125]. These complementary approaches collectively enable automated discovery of neural architectures that outperform manually designed alternatives while meeting the stringent performance and resource requirements of distributed edge-cloud deployments. For detailed technical analysis, comparative evaluation, and implementation guidelines, please see LABEL:sec:supp-resource-nas of the Supplementary Material.
@@ IV AI-Driven Resource Management
Effective resource management in ECCC systems requires intelligent orchestration of computational, communication, and storage resources across heterogeneous infrastructure to optimize system performance while meeting diverse application requirements. The section examines three complementary yet distinct aspects of AI-driven resource management. Task offloading mechanisms (Sec. IV-A) focus on intelligent decision-making for task placement across the edge-cloud continuum, considering factors such as latency requirements, energy constraints, and task dependencies to determine optimal execution locations. Resource allocation strategies (Sec. IV-B) address the dynamic distribution and management of computational resources (CPU, memory, storage) across edge nodes and cloud servers, emphasizing adaptive allocation algorithms and collaborative resource sharing approaches. Energy efficiency optimization (Sec. IV-C) concentrates on minimizing power consumption throughout the system lifecycle, including power modeling, energy-aware scheduling, and joint optimization techniques that balance performance with sustainability requirements.
@@ IV-A Task Offloading Mechanisms
TABLE IV: Summary of task offloading mechanisms, categorized by strategy type and performance optimization targets.
@@
Latency-Aware
iGATS [8] Response latency, Load balancing Outperforms baseline techniques for latency reduction High
Fan et al. [126] Task processing delay, Stability Superior performance vs. 4 comparison schemes Medium
Zhai et al. [127] Response delay, Carbon emissions 33.42× delay reduction for sensitive workloads Very High
Luckow et al. [128] Performance impact, Task placement 5-65% performance improvement across scenarios High
Energy-Efficient
PDCO [129] Energy consumption, Resource util. Approaches optimal performance in energy metrics High
Hua et al. [117] Energy consumption of mobile devices Significantly reduced energy consumption High
ETCRA [130] Energy, Time, Reliability Outperforms baselines on four key measurements High
Hu et al. [131] Energy efficiency, Service delay Achieves [O(1/V), O(V)] EE-delay tradeoff Medium
Dependency-
Aware
PRAISE [132] System benefits, Resource costs Higher benefits and lower costs vs. baselines Medium
Peng et al. [133] Computing time, Blocking probability Optimal replication with varying parameters Medium
Multi-Objective
Zhang et al. [22] Time, Cost, Load balance, Trust Found >>50% of best solutions on benchmarks High
DRLIS [134] Load balancing, Response time, Cost Up to 55% cost reduction across metrics Very High
DT-MADDPG [135] Success rate, Latency, Energy Improved success rate with reduced resource usage High
Zhang et al. [136] Time, Cost, Energy, Reliability Outperforms existing design approaches Medium
EdgeOptimizer [137] Success rate, Overall overload Improved success rate for time-critical tasks Medium
@@
Effective resource utilization and performance optimization in ECCC depend critically on sophisticated task offloading mechanisms. As depicted in Fig. 5, a framework outlines three core strategies: latency-aware, energy-efficient, and dependency-aware offloading, detailing their interactions with edge devices, cloud servers, and supporting systems to improve task placement and scheduling. Specifically, these mechanisms address distinct objectives such as real-time performance, energy consumption, and task dependencies, thereby boosting overall system efficiency. To further illustrate their practical impact, Table IV summarizes recent task offloading methods in edge-cloud environments.
Notably, the table reveals significant performance gains, such as up to a 33.42× reduction in delay for sensitive workloads [127], alongside improvements in energy consumption and system reliability, highlighting the versatility of these strategies across diverse application domains.
@@ IV-A1 Latency-Aware Offloading
Minimizing response time through latency-aware offloading represents a crucial strategy for time-sensitive applications in edge-cloud systems, with numerous approaches demonstrating significant performance gains. According to the comparative analysis in Table IV, methods like iGATS [8] and DRLIS [134] have achieved substantial reductions in processing delays through intelligent placement decisions. For instance, Lyapunov optimization techniques employed by Abouaomar et al. [138] enable dynamic resource allocation at the edge, while Poularakis et al. [18] address multidimensional constraints through joint optimization of service placement and request routing. Similarly, Fan et al. [126] propose a collaborative framework that minimizes total processing delay while ensuring queuing stability in multi-task scenarios. When considering delay-sensitive workloads specifically, Zhai et al. [127] demonstrate a remarkable 33.42× reduction in latency compared to traditional cloud deployments by intelligently distributing tasks based on their characteristics. Additionally, the double deep Q-network (DDQN) approach introduced by Pang et al. [139] effectively manages mobility while optimizing offloading decisions, while Qi et al. [140] present a two-timescale scheduling algorithm that balances service provider costs with stringent latency requirements.
@@ IV-A2 Energy-Efficient Offloading
Strategic energy management across edge-cloud systems requires balancing computation and communication costs while optimizing resource allocation through specialized techniques.
In ECCC, prolonging the battery life of edge devices necessitates balancing computation and communication energy costs, especially for battery-powered devices [131]. As shown in Table IV, numerous approaches including PDCO [129], Hua et al. [117], and ETCRA [130] have demonstrated significant improvements in energy efficiency across edge-cloud systems. Particularly for battery-powered devices, strategic decision-making has to consider both computational and communication energy costs, weighing factors like task complexity, network conditions, and hardware energy profiles [131, 141]. While offloading computationally intensive tasks to cloud resources can reduce device energy consumption, communication costs under limited bandwidth or unstable networks often counterbalance these benefits, necessitating sophisticated optimization approaches. Beyond basic consumption metrics, advanced energy-efficient mechanisms consider additional dimensions such as power control, transmission scheduling [117], and user mobility patterns [142]. Furthermore, in heterogeneous IoT deployments with diverse device capabilities and constraints, efficient resource allocation becomes increasingly crucial for sustainable service delivery [122], with some approaches even incorporating renewable energy sources to create more environmentally responsible edge-cloud ecosystems [127].
Energy-efficient offloading mechanisms strive to minimize overall energy consumption by jointly considering computation and communication costs. Additionally, optimizing power control and transmission scheduling can further enhance energy efficiency [117]. User mobility introduces another layer of complexity to energy optimization, as dynamic network conditions and varying proximity to edge servers impact energy trade-offs [142]. In heterogeneous IoT environments with diverse device capabilities and energy constraints, efficient resource allocation becomes even more crucial for energy-efficient service delivery [122]. Prioritizing workloads based on their energy requirements and leveraging renewable energy sources can contribute to a more sustainable edge-cloud ecosystem [127].
@@ IV-A3 Dependency-Aware Offloading
Complex applications often consist of multiple interconnected subtasks with dependencies, forming a dependency graph where some subtasks must finish before others can start [132]. Complex interdependence introduces complexities in task scheduling and resource allocation, potentially leading to suboptimal performance due to idle resources and increased latency if ignored. Therefore, efficient offloading mechanisms must consider these precedence constraints to optimize overall application execution. For example, the PRAISE algorithm in [132] addresses triangular dependencies by reconfiguring the task call graph into sequential layers for efficient resource allocation. Similarly, modeling task dependencies as a directed acyclic graph (DAG) optimizes offloading decisions, particularly in vehicle-edge-cloud environments [139]. Additionally, deep Q-networks (DQN) for real-time offloading in cloud-edge computing benefit from incorporating dependency awareness [143], enabling dynamic adaptation to changing environments without pre-set task priorities.
@@ IV-A4 Multi-Objective Optimization Strategies
Contemporary edge-cloud systems increasingly require sophisticated offloading strategies that simultaneously optimize multiple performance metrics, reflecting the complex trade-offs inherent in distributed computing environments. Unlike single-objective approaches that focus on isolated metrics, multi-objective optimization techniques explicitly address the conflicts between competing goals such as minimizing latency while reducing energy consumption, maximizing throughput while maintaining quality of service, and balancing computational load while ensuring system reliability [22]. Multi-objective approaches typically employ Pareto optimization methods to identify solution sets that represent optimal trade-offs between conflicting objectives, enabling system designers to select configurations that best match their specific application requirements and operational constraints [144].
Recent research demonstrates the effectiveness of multi-objective approaches across diverse scenarios. For instance, Zhang et al. [22] propose a many-objective resource scheduling model that simultaneously considers time, cost, load balance, user satisfaction, and trust factors, employing advanced evolutionary algorithms to navigate the complex solution space. Similarly, Guo et al. [145] frame service deployment and task offloading as a multi-task, high-dimensional optimization problem that incorporates network reliability and load balancing considerations alongside traditional performance metrics. Furthermore, joint optimization techniques that combine energy efficiency with latency minimization have proven particularly effective in battery-powered edge environments [117], while approaches that integrate security and privacy considerations with performance optimization address the growing demands of trust-sensitive applications [130]. The challenge lies in developing efficient algorithms that can handle the increased complexity of multi-dimensional solution spaces while maintaining real-time decision-making capabilities essential for dynamic edge-cloud environments.
@@ IV-B Resource Allocation Strategies
TABLE V: Summary of resource allocation methods.
@@
Deep Reinforcement Learning A2C-DRL [13] Edge-Cloud Environments Reward Value, Load-balancing
AC-DRL [146] 5G Edge-Cloud Continuum QoE, Energy Consumption
DA-HADRL [147] SAGINs with Cloud-Edge Energy, Latency
GRLDO [148] Client-Edge-Cloud Environment User Satisfaction
MADDPG [86] End-Edge-Cloud in 6G Networks Delay, Energy Consumption
Federated Learning FRL [93] Cloud-Edge-Terminal IoT Learning Speed, Adaptation
FDRL [95] Cloud-Edge Manufacturing Makespan, Energy Consumption
MARL-SF [149] IoV Edge Cloud Quality of Service
MARL-CA [150] Industrial Edge-Cloud Web3 Communication Rate, Stability
Optimization-Based DySCo [24] Edge-Cloud Infrastructure Deployment Cost, Delay
OTFAC [89] Cloud-Edge Collaborative IoT Delay, Energy Consumption
HDMO [136] Edge-Cloud Collaboration Time, Cost, Energy, Reliability
DaCM [151] Edge-Cloud Computing Long-term Profit, Delay
Model-Aware DSCF [152] Distributed Serverless Edge Inference Latency, Res. Consump.
DRL-LOA [153] Healthcare Cloud-Edge Power Usage, Latency
Wang et al. [154] IoT-Based Distribution Grid Convergence Rate, Stability
Domain-Specific Das et al. [155] Drone Communication Networks Latency, Scalability, Resource Util.
DA-DDQN [139] Vehicle-Edge-Cloud Latency, Energy Consumption
QRSRL [156] Train-Edge-Cloud Smart Systems Task Processing Delay
Informer-A3C [157] Vehicular Edge Cloud Computing Content Access Cost
@@
Strategic allocation of computational resources across the edge-cloud continuum remains fundamental for maximizing performance in AI-driven collaborative environments. Accordingly, we explore three pivotal strategies: dynamic resource allocation, collaborative resource management, and reinforcement learning-based optimization, each designed to enhance task execution and system efficiency. By effectively distributing resources, advanced allocation approaches address the challenges of varying workloads and resource constraints in distributed environments. To illustrate their practical impact, Table V provides a comparative analysis of recent dynamic resource allocation methods in ECCC systems.
Notably, the table reveals a wide range of application domains, from IoT networks to vehicular edge clouds, alongside key performance metrics like latency, energy consumption, and quality of service, highlighting the adaptability and effectiveness of these strategies in diverse scenarios.
Resource allocation strategies in ECCC systems can be systematically organized according to their adaptation mechanisms and coordination approaches: (1) Temporal adaptation strategies (dynamic resource allocation) that respond to time-varying workloads and resource availability through real-time monitoring and reactive adjustment mechanisms; (2) Spatial coordination strategies (collaborative resource management) that optimize resource distribution across geographically distributed edge nodes and cloud infrastructure through inter-node cooperation and hierarchical coordination; (3) Learning-based optimization strategies (reinforcement learning approaches) that continuously improve resource allocation decisions through environmental feedback and experience accumulation. The classification framework emphasizes the complementary nature of these approaches, where temporal adaptation handles immediate demand fluctuations, spatial coordination optimizes cross-infrastructure efficiency, and learning-based methods provide long-term performance improvement through adaptive policy refinement.

### v5

@@ III-C Resource-Aware Model Optimization
Resource-aware model optimization focuses on dynamically adapting AI models to the heterogeneous and constrained computational environments found in ECCC systems. Building upon foundational techniques discussed in prior surveys [qu2025mobile], we classify advanced resource-aware strategies into three primary approaches based on their operational focus. Dynamic model selection adapts to resource availability by choosing the most appropriate model from a pool of candidates at runtime. Model splitting and offloading partitions model components across the edge-cloud infrastructure to balance latency and computational load. Reinforcement learning-based optimization uses learning-based methods for adaptive resource allocation and decision-making. Working together, these strategies collectively address the challenges of resource heterogeneity and dynamic operational constraints. For instance, in AR applications, model splitting can offload scene recognition to an edge server while keeping real-time pose rendering on the local device to ensure a fluid user experience [hua2024energyefficient]. Similarly, reinforcement learning can optimize resource allocation in IoT monitoring applications to improve stability and reduce failure rates [xu2024fusion, goudarzi2024mmddrl]. For comprehensive technical details on each approach, please see LABEL:sec:supp-resource-aware of the Supplementary Material.
@@ III-D Neural Architecture Search
NAS automates the design of efficient neural networks, a critical task for creating models tailored to the specific constraints of ECCC. In line with established surveys, we categorize NAS techniques based on their underlying search strategy. Building upon this foundation, the three dominant paradigms are: reinforcement learning-based search, where an agent learns optimal architectures through environmental feedback; evolutionary algorithm-based search, which mimics natural selection to evolve populations of architectures; and gradient-based search, which uses gradient descent for efficient optimization of a continuous architectural representation [goudarzi2023distributed, he2024efficient, zhang2024manyobjective]. Each of these paradigms offers distinct advantages for discovering architectures that balance performance with the resource limitations of edge devices. For a detailed technical analysis of each search strategy, please see LABEL:sec:supp-resource-nas of the Supplementary Material.
III-E Lessons Learned on Model Optimization
Our analysis of model optimization techniques for ECCC systems reveals several critical insights. A recurring theme is the inherent trade-off between performance and efficiency, which manifests differently across various optimization strategies. Key lessons learned include: (1) Model compression techniques, such as pruning, quantization, and knowledge distillation, effectively reduce model complexity but often incur accuracy degradation, with the optimal balance being highly model- and task-dependent. (2) Model adaptation methods like federated and continual learning address data privacy and model evolution but introduce significant communication bottlenecks and are sensitive to non-IID data distributions, which remains a primary research obstacle. (3) Neural architecture search automates the design of efficient models, yet its substantial computational overhead limits direct application in resource-constrained edge scenarios, thereby motivating the development of lightweight and one-shot NAS methods. (4) A holistic optimization strategy must dynamically consider device capabilities, network conditions, and application requirements, where dynamic model selection and partitioning are crucial for achieving resource-aware optimization. Notably, the model optimization techniques detailed in this section primarily address “design-time” challenges, creating AI models that are inherently smaller, faster, and more adaptable. Accordingly, an efficiently optimized model fundamentally simplifies the subsequent “run-time” challenges of task offloading and resource allocation, whereby the AI-driven resource management strategies discussed in Sec. IV are specifically designed to effectively deploy and run the AI models optimized in this section, thus serving as a critical foundation for translating model-level improvements into system-wide performance gains.
@@ IV AI-Driven Resource Management
Resource Management (§IV) Task Offloading (§IV-A) Resource Allocation (§IV-B) Energy Efficiency (§IV-C) Latency-Aware Energy-Efficient Dependency-Aware Multi-Objective Dynamic Allocation Collaborative Management RL-based Allocation Power Management Energy-Aware Scheduling Green Computing [sun2025latencyaware, wang2024deep, zhai2024edgecloud] [su2024primaldualbased, hua2024energyefficient, li2024energyaware] [feng2024dependencyaware, pang2024multimobile, chen2024realtime] [zhang2024manyobjective, jain2022latencymemory, guo2023edge] [farhoudi2023qosaware, kim2024collaborative, chen2024dynamica] [gan2024optimal, wang2024cloudedgeend, moraes2019pub] [binh2024reinforcement, huang2024joint, afachao2024efficient] [georgiou2021comprehensive, yao2023edgecloud, zhang2024dvfo] [nieto2024deep, zhang2024multitargetaware, bensalah2024vnf] [arroba2024sustainable, chen2024fast, abouaomar2021resource]
Figure 6: Taxonomy of AI-driven resource management.
Building upon the foundation of optimized AI models from Sec. III, this section investigates the critical challenge of “run-time” resource management. Specifically, while model optimization prepares AI models for efficient execution, resource management strategies govern their actual deployment, scheduling, and orchestration within the dynamic and heterogeneous ECCC environment. To provide a structured overview of this domain, Fig. 6 presents a hierarchical taxonomy of AI-driven resource management, organized into three primary branches. First, task offloading mechanisms (Sec. IV-A) determine optimal task placement. Second, resource allocation strategies (Sec. IV-B) manage the distribution of computational resources. Third, energy efficiency optimization (Sec. IV-C) focuses on minimizing power consumption. Accordingly, the following subsections are structured according to this taxonomy, exploring how approaches within each category enable the full potential of optimized models to be realized, thereby ensuring robust and efficient system performance.
@@ IV-A Task Offloading Mechanisms
TABLE V: Summary of task offloading mechanisms, categorized by strategy type and performance optimization targets.
@@
Latency-Aware iGATS [sun2025latencyaware] Response latency, Load balancing Outperforms baseline techniques for latency reduction High
Fan et al. [fan2024collaborative] Task processing delay, Stability Superior performance vs. 4 comparison schemes Medium
Zhai et al. [zhai2024edgecloud] Response delay, Carbon emissions 33.42× delay reduction for sensitive workloads Very High
Luckow et al. [luckow2021exploring] Performance impact, Task placement 5-65% performance improvement across scenarios High
Energy-Efficient PDCO [su2024primaldualbased] Energy consumption, Resource util. Approaches optimal performance in energy metrics High
Hua et al. [hua2024energyefficient] Energy consumption of mobile devices Significantly reduced energy consumption High
ETCRA [li2024energyaware] Energy, Time, Reliability Outperforms baselines on four key measurements High
Hu et al. [hu2022energy] Energy efficiency, Service delay Achieves [O(1/V), O(V)] EE-delay tradeoff Medium
Dependency- Aware PRAISE [feng2024dependencyaware] System benefits, Resource costs Higher benefits and lower costs vs. baselines Medium
Peng et al. [peng2024redundancy] Computing time, Blocking probability Optimal replication with varying parameters Medium
Multi-Objective Zhang et al. [zhang2024manyobjective] Time, Cost, Load balance, Trust Found >>50% of best solutions on benchmarks High
DRLIS [wang2024deep] Load balancing, Response time, Cost Up to 55% cost reduction across metrics Very High
DT-MADDPG [wang2024cooperative] Success rate, Latency, Energy Improved success rate with reduced resource usage High
Zhang et al. [zhang2024hybrid] Time, Cost, Energy, Reliability Outperforms existing design approaches Medium
EdgeOptimizer [qiao2024edgeoptimizer] Success rate, Overall overload Improved success rate for time-critical tasks Medium
@@
Effective resource utilization and performance optimization in ECCC depend critically on sophisticated task offloading mechanisms. As depicted in Fig. 7, a framework outlines three core strategies: latency-aware, energy-efficient, and dependency-aware offloading, detailing their interactions with edge devices, cloud servers, and supporting systems to improve task placement and scheduling. Specifically, these mechanisms address distinct objectives such as real-time performance, energy consumption, and task dependencies, thereby boosting overall system efficiency. To further illustrate their practical impact, Table V summarizes recent task offloading methods in edge-cloud environments. Notably, the table reveals significant performance gains, such as up to a 33.42× reduction in delay for sensitive workloads [zhai2024edgecloud], alongside improvements in energy consumption and system reliability, highlighting the versatility of these strategies across diverse application domains.
@@ IV-A1 Latency-Aware Offloading
Minimizing response time through latency-aware offloading represents a crucial strategy for time-sensitive applications in edge-cloud systems, with numerous approaches demonstrating significant performance gains. According to the comparative analysis in Table V, methods like iGATS [sun2025latencyaware] and DRLIS [wang2024deep] have achieved substantial reductions in processing delays through intelligent placement decisions. For instance, Lyapunov optimization techniques employed by Abouaomar et al. [abouaomar2021resource] enable dynamic resource allocation at the edge, while Poularakis et al. [poularakis2019joint] address multidimensional constraints through joint optimization of service placement and request routing. Similarly, Fan et al. [fan2024collaborative] propose a collaborative framework that minimizes total processing delay while ensuring queuing stability in multi-task scenarios. When considering delay-sensitive workloads specifically, Zhai et al. [zhai2024edgecloud] demonstrate a remarkable 33.42× reduction in latency compared to traditional cloud deployments by intelligently distributing tasks based on their characteristics. Additionally, the double deep Q-network (DDQN) approach introduced by Pang et al. [pang2024multimobile] effectively manages mobility while optimizing offloading decisions, while Qi et al. [qi2024service] present a two-timescale scheduling algorithm that balances service provider costs with stringent latency requirements. Furthermore, the local computation delay of a task is not static; it can be significantly reduced through the model compression techniques detailed in Sec. III-A, such as pruning and quantization, thereby making local execution more feasible.
@@ IV-A2 Energy-Efficient Offloading
Strategic energy management across edge-cloud systems requires balancing computation and communication costs while optimizing resource allocation through specialized techniques. For instance, in mobile AR, offloading computationally intensive rendering tasks from a headset to a nearby edge server can significantly conserve battery life, extending the device’s operational time while maintaining an interactive experience. In ECCC, prolonging the battery life of edge devices necessitates balancing computation and communication energy costs, especially for battery-powered devices [hu2022energy]. As shown in Table V, numerous approaches including PDCO [su2024primaldualbased], Hua et al. [hua2024energyefficient], and ETCRA [li2024energyaware] have demonstrated significant improvements in energy efficiency across edge-cloud systems. Particularly for battery-powered devices, strategic decision-making has to consider both computational and communication energy costs, weighing factors like task complexity, network conditions, and hardware energy profiles [hu2022energy, he2019collaborative]. Notably, the choice of model architecture has a decisive impact on computational energy consumption, a key optimization target for the neural architecture search (NAS) techniques discussed in Sec. III-D. While offloading computationally intensive tasks to cloud resources can reduce device energy consumption, communication costs under limited bandwidth or unstable networks often counterbalance these benefits, necessitating sophisticated optimization approaches. Beyond basic consumption metrics, advanced energy-efficient mechanisms consider additional dimensions such as power control, transmission scheduling [hua2024energyefficient], and user mobility patterns [arroba2024sustainable]. Furthermore, in heterogeneous IoT deployments with diverse device capabilities and constraints, efficient resource allocation becomes increasingly crucial for sustainable delivery [chen2024fast], with some approaches even incorporating renewable energy sources to create more environmentally responsible edge-cloud ecosystems [zhai2024edgecloud].
Energy-efficient offloading mechanisms strive to minimize overall energy consumption by jointly considering computation and communication costs. Additionally, optimizing power control and transmission scheduling can further enhance energy efficiency [hua2024energyefficient]. User mobility introduces another layer of complexity to energy optimization, as dynamic network conditions and varying proximity to edge servers impact energy trade-offs [arroba2024sustainable]. In heterogeneous IoT environments with diverse device capabilities and energy constraints, efficient resource allocation becomes even more crucial for energy-efficient service delivery [chen2024fast]. Prioritizing workloads based on their energy requirements and leveraging renewable energy sources can contribute to a more sustainable edge-cloud ecosystem [zhai2024edgecloud].
@@ IV-A3 Dependency-Aware Offloading
Complex applications often consist of multiple interconnected subtasks with dependencies, forming a dependency graph where some subtasks must finish before others can start [feng2024dependencyaware]. Complex interdependence introduces complexities in task scheduling and resource allocation, potentially leading to suboptimal performance due to idle resources and increased latency if ignored. Therefore, efficient offloading mechanisms must consider these precedence constraints to optimize overall application execution. For example, the PRAISE algorithm in [feng2024dependencyaware] addresses triangular dependencies by reconfiguring the task call graph into sequential layers for efficient resource allocation. Similarly, modeling task dependencies as a directed acyclic graph (DAG) optimizes offloading decisions, particularly in vehicle-edge-cloud environments [pang2024multimobile]. Additionally, deep Q-networks (DQN) for real-time offloading in cloud-edge computing benefit from incorporating dependency awareness [chen2024realtime], enabling dynamic adaptation to changing environments without pre-set task priorities.
@@ IV-A4 Multi-Objective Optimization Strategies
Contemporary edge-cloud systems increasingly demand sophisticated offloading strategies that simultaneously optimize conflicting performance metrics like latency, energy consumption, and reliability through Pareto optimization methods [zhang2024manyobjective, jain2022latencymemory]. Approaches like the many-objective resource scheduling model proposed by Zhang et al. [zhang2024manyobjective] effectively navigate complex solution spaces considering time, cost, and load balance, whilst other frameworks frame service deployment as high-dimensional problems incorporating network reliability [guo2023edge]. Additionally, joint optimization techniques combining energy efficiency with latency minimization prove effective in battery-powered environments [hua2024energyefficient], and integrating security considerations addresses trust-sensitive application demands [li2024energyaware], highlighting the critical need for efficient algorithms capable of handling multi-dimensional constraints in real-time.
@@ IV-B Resource Allocation Strategies
TABLE VI: Summary of resource allocation methods.
@@
Deep Reinforcement Learning A2C-DRL [lu2024a2cdrl] Edge-Cloud Environments Reward Value, Load-balancing
AC-DRL [nieto2024deep] 5G Edge-Cloud Continuum QoE, Energy Consumption
DA-HADRL [huang2024joint] SAGINs with Cloud-Edge Energy, Latency
GRLDO [li2024offloading] Client-Edge-Cloud Environment User Satisfaction
MADDPG [she2024efficient] End-Edge-Cloud in 6G Networks Delay, Energy Consumption
Federated Learning FRL [kim2024collaborative] Cloud-Edge-Terminal IoT Learning Speed, Adaptation
FDRL [wang2024federated] Cloud-Edge Manufacturing Makespan, Energy Consumption
MARL-SF [lu2024dynamic] IoV Edge Cloud Quality of Service
MARL-CA [yang2024multiagent] Industrial Edge-Cloud Web3 Communication Rate, Stability
Optimization-Based DySCo [bensalah2024vnf] Edge-Cloud Infrastructure Deployment Cost, Delay
OTFAC [gan2024optimal] Cloud-Edge Collaborative IoT Delay, Energy Consumption
HDMO [zhang2024hybrid] Edge-Cloud Collaboration Time, Cost, Energy, Reliability
DaCM [wang2024efficient] Edge-Cloud Computing Long-term Profit, Delay
Model-Aware DSCF [ko2024dynamic] Distributed Serverless Edge Inference Latency, Res. Consump.
DRL-LOA [saranya2024enhanced] Healthcare Cloud-Edge Power Usage, Latency
Wang et al. [wang2024cloudedgeend] IoT-Based Distribution Grid Convergence Rate, Stability
Domain-Specific Das et al. [das2024edgecloud] Drone Communication Networks Latency, Scalability, Resource Util.
DA-DDQN [pang2024multimobile] Vehicle-Edge-Cloud Latency, Energy Consumption
QRSRL [zhu2024collaborative] Train-Edge-Cloud Smart Systems Task Processing Delay
Informer-A3C [zhu2024vehicular] Vehicular Edge Cloud Computing Content Access Cost
@@
Strategic allocation of computational resources across the edge-cloud continuum remains fundamental for maximizing performance in AI-driven collaborative environments. Accordingly, we explore three pivotal strategies: dynamic resource allocation, collaborative resource management, and reinforcement learning-based optimization, each designed to enhance task execution and system efficiency. By effectively distributing resources, advanced allocation approaches address the challenges of varying workloads and resource constraints in distributed environments. To illustrate their practical impact, Table VI provides a comparative analysis of recent dynamic resource allocation methods in ECCC systems. Notably, the table reveals a wide range of application domains, from IoT networks to vehicular edge clouds, alongside key performance metrics like latency, energy consumption, and quality of service, highlighting the adaptability and effectiveness of these strategies in diverse scenarios.
Resource allocation strategies in ECCC systems can be systematically organized according to their adaptation mechanisms and coordination approaches: (1) Temporal adaptation strategies (dynamic resource allocation) that respond to time-varying workloads and resource availability through real-time monitoring and reactive adjustment mechanisms; (2) Spatial coordination strategies (collaborative resource management) that optimize resource distribution across geographically distributed edge nodes and cloud infrastructure through inter-node cooperation and hierarchical coordination; (3) Learning-based optimization strategies (reinforcement learning approaches) that continuously improve resource allocation decisions through environmental feedback and experience accumulation. Accordingly, the classification framework emphasizes the complementary nature of these approaches, where temporal adaptation handles immediate demand fluctuations, spatial coordination optimizes cross-infrastructure efficiency, and learning-based methods provide long-term performance improvement through adaptive policy refinement.